OpenStudy (czesc):

The components of a position vector of a particle moving in the plane are . What is the distance traveled by the particle from t = 0 to t = π?

OpenStudy (anonymous):

Do you know the arclength formula?

OpenStudy (czesc):

@wio @raffle_snaffle now that I know the formula, how do I apply it here? thanks for the responses guys!

OpenStudy (anonymous):

Given the position vector \(\mathbf p(t)\), the arc lenth between \(t_1\) and \(t_2\) is: \[ L=\int\limits_{t_1}^{t_2} \|\mathbf p'(t)\|~dt \]Given that the velocity vector \(\mathbf v(t)\) is equal to \(\mathbf p'(t)\) we can write: \[ L=\int\limits_{t_1}^{t_2} \|\mathbf v(t)\|~dt \]

OpenStudy (anonymous):

`from t = 0 to t = π?` This means \(t_1=0\) and \(t_2=\pi\).

OpenStudy (anonymous):

So plugging it all in, we have: \[ L = \int_0^\pi \|\langle t^2, \sin(t)\rangle\|~dt \]

OpenStudy (anonymous):

Do you think you can do the rest?

OpenStudy (czesc):

@wio , Thanks for all of that, I understand a lot more now, but how do I integrate the vector components? (sorry if I sound arrogant)

OpenStudy (anonymous):

There is no vector, because we are taking the magnitude of the vector.

OpenStudy (czesc):

@wio So how do I integrate <t^2, sin(t)>dt from 0 to pi

OpenStudy (anonymous):

Do you know how to find the magnitude of a vector?

OpenStudy (anonymous):

\[ \|\langle t^2,\sin(t)\rangle\| = \sqrt{\langle t^2,\sin(t)\rangle\cdot \langle t^2,\sin(t)\rangle} = \sqrt{t^4+\sin^2(t)} \]

OpenStudy (czesc):

Oh so it would be sqrt(v(x)^2+v(y)^2) then it would just be: \[\int\limits_{0}^{\pi} \sqrt{t^4+\sin^2(t)}=10.8265\] correct?

OpenStudy (czesc):

@wio Thank you very much for being so detailed and patient with me I highly appreciate it.

OpenStudy (anonymous):

Yes